OpenStudy (anonymous):

x^4 -81/x+3 Explain please!

OpenStudy (anonymous):

\[x^4 -81 \] There is some way you can expand this in the form of (x+3)

OpenStudy (anonymous):

I just can't figure it out without getting the -81

OpenStudy (anonymous):

I always get positive

OpenStudy (campbell_st):

factorise the numerator \[\frac{(x^2 - 9)(x^+9)}{x+3}\] (x^2 - 9) is the difference of 2 squares and can be factorised \[\frac{(x -3)(x+3)(x^2 +9)}{(x+3)}\] now you can eliminate the common factor ( x+3) \[(x -3)(x^2 + 9)\]

OpenStudy (campbell_st):

oops 1s\[\frac{(x^2 -9)(x^2 +9)}{(x +3)}\]t factorisation should be